OpenStudy (anonymous):

the lim as x approaches positive infinity of (x-(1/2x))/(2x+6x) is?

OpenStudy (anonymous):

how do you solve for an infinite limit?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}\frac{x-\frac{1}{2}x}{2x+6x}=\lim_{x \rightarrow \infty}\frac{\frac{1}{2}x}{8x}=\lim_{x \rightarrow \infty}\frac{\frac{1}{2}}{8}=\frac{1}{16}\]

OpenStudy (anonymous):

very easy you should just consider the high rate of varaiable in nominator and denominator so answer is (1/2x)/8x=1/16

OpenStudy (anonymous):

Okay so... I messed up the problem. Here it is \[\lim_{x \rightarrow \infty}\frac{ x- \frac{ 1 }{ 2x } }{ 2x+ \frac{ 1 }{ 6x } }\]

OpenStudy (anonymous):

HAHA xD

OpenStudy (anonymous):

I was wondering why your work was so much simpler than what I thought, I realized I messed up.

OpenStudy (anonymous):

That's why everbody has to use \[\mathrm{LaTeX}\] The answer is 1/2 and the And the procediment is similar.

OpenStudy (anonymous):

I've got it now! Thank you!